### v4

@@ 8.1 Universal model
We will discuss the model appropriate for the cuprates here, although the results are ‘universal’ as they apply also to other metallic quantum phase transitions in the presence of spatial disorder [90] which are relevant to numerous other correlated electron materials. Our model for the hole-doped cuprates is a FL*-FL quantum phase transition (see Fig. 37) without symmetry breaking, and in the clean limit this is distinct from metallic quantum phase transitions with symmetry breaking. The latter are further distinguished in the clean limit by cases in which the order parameters reside at zero or non-zero momenta [232] which have dynamic critical exponents $z=3$ and $z=2$ respectively. The remarkable fact is that spatial disorder unifies these 3 classes into essentially the same universal theory [90] which we summarize here. However, we note that there is an enhancement of thermopower which is special to the FL*-FL transition, as discussed elsewhere [109], and this compares well with experiments on the cuprates [110, 111, 112].
@@
Figure 36: On the left is the cuprate phase diagram from Ref. [2]; annotations in blue have been added. On the right is the ALM, with the Higgs fields $\Phi$ and $B$. In Section 8 we consider the transition from the FL* pseudogap to the Fermi liquid focusing on the Higgs field $\Phi$, while setting $B=0$. The transition from the pseudogap to the d-SC was discussed in Section 6 as a theory of the dynamics of $B$, while setting $\Phi$ to a non-zero constant which determined the magnitude of the pseudogap. See the Appendix for the complete gauge theory of $B$ and $\Phi$ for both transitions.
@@
In Section 6 we have considered the low temperature fate of the FL* pseudogap associated with condensation of the Higgs field $B$. In this analysis, we treated $\Phi$ simply as a c-number constant. In the present section, $\Phi$ and $B$ will exchange roles. In considering the transition from FL* to FL in Fig. 8 and Fig. 36, we consider the dynamics of the $\Phi$ field, and the fermions in the top two layers in Fig. 21. In FL*, we can safely set $B=0$ (apart for some thermal fluctuations which were needed in Section 5.1). In the FL phase of Fig. 8, we see that the $f_{1}$ fermions of the middle layer form a trivial rung-singlet state with the $f$ fermions in the bottom layer: this confinement of $f_{1}$ and $f$ is not associated with $B$, but by the confinement of a $\widetilde{\rm SU}(2)$ gauge field [48, 91, 152] (see the Appendix). Consequently, it is safe to ignore $B$ in the quantum criticality of the FL* to FL transition [48, 91, 233]. So we will only consider the $\Phi$ boson here, along with the $c_{\alpha}$ and $f_{1}$ fermions in the top two layers of the ALM.
@@
We therefore consider the $T>0$ quantum criticality of the FL* to FL transition sketched in Fig. 37, associated with the Higgs transition in $\Phi$.
@@
Figure 37: Schematic phase diagram of a single-band metal with a Fermi surface volume changing quantum transition without a broken symmetry on either side of the transition. For the Kondo lattice, the transition is ‘inverted’ as shown in Fig. 9, with the Higgs field $\Phi$ condensed on the FL side. The ‘naked’ quantum critical point at $s_{c}$ is the same as that at $p_{c}$ in Fig. 1. This quantum critical point is expected to the unstable to a superconducting dome [48, 91], with the superconducting state smoothly connected across $s_{c}$, as shown in Fig. 1 at $p_{c}$. This is an important difference from quantum phase transitions in metals with symmetry breaking, for which the quantum phase transition is present inside the superconductor.
@@
We could study this transition using the ‘Kondo breakdown’ dynamic mean field theories [92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106], which map the transition to SYK-type Green’s functions studied in Section 7. This could apply in very clean Kondo lattice systems over a significant intermediate temperature range, but does have the serious issue of a non-vanishing entropy density in the zero temperature limit [92]. Here, we will instead treat the spatial dimensionality seriously, and focus on the low energy, long-wavelength structure. Spatially random couplings play a crucial role in such an analysis, and lead to new Griffiths effects associated with the ‘foot’ of strange metal behavior [116, 117, 114] which we will discuss in Section 9.
@@
At the mean-field level the FL* to FL transition is described by $\mathcal{H}_{\rm KLmf}$ in Eq. (2.4). We now include spatial dependence in all fields, allow for quenched spatial disorder, and obtain the (universal) Lagrangian of the 2D-YSYK model
@@
We are using the convention that quenched random variables (with no dynamics and a fixed, random dependence on $\bm{r}$}) are colored in purple. For the Kondo lattice case, there should also be a U(1) gauge field associated with Eq. (2.5) in the Lagrangian in Eq. (8.6), but we have dropped it as it does not significantly modify the critical behavior [234, 217]. For the single-band case, there is a $\widetilde{\rm SU}(2)\times{\rm U}(1)$ gauge field [48, 91, 233], as we review in the Appendix (the tilde is to distinguish from the distinct SU(2) gauge field of Sections 3-6); the U(1) gauge field is the local constraint in the ${\bm{S}}_{1}$ layer (as in Eq. (2.5)), and the $\widetilde{\rm SU}(2)$ gauge field is the singlet projection between the ${\bm{S}}_{1}$ and ${\bm{S}}_{2}$ layers, and we ignore these gauge fields for similar reasons. The dispersions $\varepsilon({\bm{k}})$ and $\varepsilon_{1}({\bm{k}})$ arise from the $t_{{\bm{i}}{\bm{j}}}$ and $t_{1{\bm{i}}{\bm{j}}}$ in Eq. (2.4), $s$ is the tuning parameter in Fig. 37, and we have added various in an effective action for $\Phi$.
@@ 8.3 With spatial randomness
The absence of the $\omega^{-2/3}$ term in Eq. (8.12) is a strong indication that the resolution of the strange metal problem cannot come from a clean system. There can be umklapp processes which dissipate momentum, but these require special features of the Fermi surface to survive at low momentum. A theory with only potential scattering disorder as in Eq. (8.7), i.e. $g\neq 0$, $v\neq 0$, but $g^{\prime}=0$, is also not sufficient [234, 244]: it leads to marginal Fermi-liquid behavior in the electron self energy, but no strange metal behavior in transport. So for a generic and universal theory of strange metals, the influence of disorder with $g$, $g^{\prime}$, and $v$ all non-zero should be considered. The solution of Eq. (8.10) yields a boson Green’s function which has a diffusive form at the critical point [245]
@@
This expression shows that the residual resistivity $\rho_{0}$ at $T=0$ is determined by the elastic scattering rate $1/\tau_{e}\sim{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}v^{2}}$, as in a disordered Fermi liquid. The inelastic processes lead to a frequency and temperature dependence which matches precisely with the observational form in Eq. (8.2). An important feature is that of the two processes contributing to the electron inelastic scattering rate $1/\tau_{\rm in}$ in Eq. (8.15), only one contributes to the inelastic transport rate $1/\tau_{\rm trans}$. The processes involving the spatially uniform interaction $g$ and the impurity potential $v$ cancel out in the computation of the conductivity from Fig. 38, and only those involving the spatially random interaction $g^{\prime}$ survive [90]. A consequence of this cancellation is that the constant $\alpha$ in Eq. (8.5) approaches $\alpha=\pi/2$ for the quasiparticle $m^{\ast}$ in the limit ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}g^{\prime}}\gg g$ [235], and decreases from this value as $g$ is increased [90].
To summarize, the conductivity of the theory $\mathcal{L}_{\Phi}$ yields the strange metal conductivity in Eq. (8.1), with $\rho_{0}\sim{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}v^{2}}$ and $A\sim{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}g^{\prime 2}}$. Note that the value of $g$ does not make a direct difference to the value of the linear-$T$ resistivity, although it does affect the marginal Fermi liquid behavior of the electron self energy, as noted in Eq. (8.15). It is also notable that the residual resistivity and linear-$T$ resistivity slope are determined by different sources of disorder: those in Eqs. (8.7) and (8.8) respectively. This distinction should be important in understanding trends in observations [246, 247].
@@
We note that a full numerical solution of Eq. (8.10) at $g=0$ has been presented by Li et al. [108], including results for the conductivity and the onset of superconductivity. Their results for the transport scattering time $\tau_{\rm trans}$ are shown in Fig. 39, along with the corresponding optical conductivity observations of Michon et al [227]. Both obey the Planckian scaling form in Eq. (8.2) reasonably well.
@@ 9.5 Photoemission
This research was supported by NSF Grant DMR-2245246 and by the Simons Collaboration on Ultra-Quantum Matter which is a grant from the Simons Foundation (651440, S. S.). The Flatiron Institute is a division of the Simons Foundation.
Appendix. Complete gauge theory of the Ancilla Layer Model

### v6

@@ 8.1 Universal model
We will discuss the model appropriate for the cuprates here, although the results can be extended to other metallic quantum phase transitions with or without disorder. As we have noted in Section 1.2, in the clean case when umklapp scattering is suppressed, although there can be no quasiparticles around the Fermi surface, the transport is nevertheless Fermi liquid-like: this will be described in Section 8.2. So we will focus on the disordered case, with the most important disorder being of the Harris type, which produces local variations in the value of the coupling tuning the system across the quantum critical point.
Our model for the hole-doped cuprates is a FL*-FL quantum phase transition (see Fig. 37) without symmetry breaking, and in the clean limit this is distinct from metallic quantum phase transitions with symmetry breaking. The latter are further distinguished in the clean limit by cases in which the order parameters reside at zero or non-zero momenta [14] which have dynamic critical exponents $z=3$ and $z=2$ respectively. The remarkable fact is that spatial disorder unifies these 3 classes into essentially the same universal theory [122] which we summarize below. However, we note that there is an enhancement of thermopower which is special to the FL*-FL transition, as discussed elsewhere [143], and this compares well with experiments on the cuprates [144, 145, 146].
@@
Figure 36: On the left is the cuprate phase diagram from Ref. [5]; annotations in blue have been added. On the right is the ALM, with the Higgs fields $\Phi$ and $B$. In Section 8 we consider the transition from the FL* pseudogap to the Fermi liquid focusing on the Higgs field $\Phi$, while setting $B=0$. The transition from the pseudogap to the d-SC was discussed in Section 6 as a theory of the dynamics of $B$, while setting $\Phi$ to a non-zero constant which determined the magnitude of the pseudogap. See Appendix B for the complete gauge theory of $B$ and $\Phi$ for both transitions.
@@
In Section 6 we have considered the low temperature fate of the FL* pseudogap associated with condensation of the Higgs field $B$. In this analysis, we treated $\Phi$ simply as a c-number constant. In the present section, $\Phi$ and $B$ will exchange roles. In considering the transition from FL* to FL in Fig. 8 and Fig. 36, we consider the dynamics of the $\Phi$ field, and the fermions in the top two layers in Fig. 21. In FL*, we can safely set $B=0$ (apart for some thermal fluctuations which were needed in Section 5.1). In the FL phase of Fig. 8, we see that the $f_{1}$ fermions of the middle layer form a trivial rung-singlet state with the $f$ fermions in the bottom layer: this confinement of $f_{1}$ and $f$ is not associated with $B$, but by the confinement of a $\widetilde{\rm SU}(2)$ gauge field [4, 123, 191] (see Appendix B). Consequently, it is safe to ignore $B$ in the quantum criticality of the FL* to FL transition [4, 123, 278]. So we will only consider the $\Phi$ boson here, along with the $c_{\alpha}$ and $f_{1}$ fermions in the top two layers of the ALM.
@@
We therefore consider the $T>0$ quantum criticality of the FL* to FL transition [2, 3] sketched in Fig. 37, associated with the Higgs transition in $\Phi$ [4].
@@
Figure 37: Schematic phase diagram of a single-band metal with a Fermi surface volume changing quantum transition without a broken symmetry on either side of the transition. For the Kondo lattice, the transition is ‘inverted’ as shown in Fig. 9, with the Higgs field $\Phi$ condensed on the FL side. The ‘naked’ quantum critical point at $s_{c}$ is the same as that at $p_{c}$ in Fig. 1. This quantum critical point is expected to the unstable to a superconducting dome [4, 123], with the superconducting state smoothly connected across $s_{c}$, as shown in Fig. 1 at $p_{c}$. This is an important difference from quantum phase transitions in metals with symmetry breaking, for which the quantum phase transition is also present inside the superconductor.
@@
Dynamic mean-field theories. We could study this transition using the dynamic mean field theories [119, 124, 125, 126, 15, 16, 127, 128, 129, 130, 131, 132, 17, 133, 134, 135, 136, 137], which map the transition to self-consistent Green’s functions similar to those studied in Section 7. Such theories are in the strong umklapp limit, and could apply in very clean Kondo lattice systems over a significant intermediate temperature range. However, we expect umklapp to be suppressed at low temperatures, and the emergence of a conserved momentum leads to absence of strange metal behavior, as we discuss in Section 8.2; similar comments apply to other numerical studies [279, 280]. Moreover, the dynamic mean-field theories also have the serious issue of a non-vanishing entropy density in the zero temperature limit [119].
Here, we will instead treat the spatial dimensionality seriously, and focus on the low energy, long-wavelength structure. Spatially random couplings play a crucial role in such an analysis, and lead to new Griffiths effects associated with the ‘foot’ of strange metal behavior [149, 150, 121] which we will discuss in Section 9.
@@
At the mean-field level the FL* to FL transition is described by $\mathcal{H}_{\rm KLmf}$ in Eq. (2.4). We now include spatial dependence in all fields, allow for quenched spatial disorder, and obtain the (universal) Lagrangian of the two-dimensional Yukawa-Sadhdev-Ye-Kitaev (2D-YSYK) model
@@
We are using the convention that quenched random variables (with no dynamics and a fixed, random dependence on $\bm{r}$}) are colored in purple. For the Kondo lattice case, there should also be a U(1) gauge field associated with Eq. (2.5) in the Lagrangian in Eq. (8.6), but we have dropped it as it does not significantly modify the critical behavior [281, 260]. For the single-band case, there is a $\widetilde{\rm SU}(2)\times{\rm U}(1)$ gauge field [4, 123, 278], as we review in Appendix B (the tilde is to distinguish from the distinct SU(2) gauge field of Sections 3-6); the U(1) gauge field is the local constraint in the ${\bm{S}}_{1}$ layer (as in Eq. (2.5)), and the $\widetilde{\rm SU}(2)$ gauge field is the singlet projection between the ${\bm{S}}_{1}$ and ${\bm{S}}_{2}$ layers, and we ignore these gauge fields for similar reasons. The dispersions $\varepsilon({\bm{k}})$ and $\varepsilon_{1}({\bm{k}})$ arise from the $t_{{\bm{i}}{\bm{j}}}$ and $t_{1{\bm{i}}{\bm{j}}}$ in Eq. (2.4), $s$ is the tuning parameter in Fig. 37, and we have added various in an effective action for $\Phi$.
@@ 8.3 With spatial randomness
The absence of the $\omega^{-2/3}$ term in Eq. (8.12) is a strong indication that the resolution of the strange metal problem cannot come from a clean system. There can be umklapp processes which dissipate momentum, but these require special features of the Fermi surface to survive at low momentum. Recent nonlinear THz 2D coherent spectroscopy observations on the cuprates [291] observe a rapid momentum relaxation rate, supporting models in which disorder plays a central role.
A theory with only potential scattering disorder as in Eq. (8.7), i.e. $g\neq 0$, $v\neq 0$, but $g^{\prime}=0$, is also not sufficient [281, 292]: it leads to marginal Fermi-liquid behavior in the electron self energy, but no strange metal behavior in transport. So for a generic and universal theory of strange metals, the influence of disorder with $g$, $g^{\prime}$, and $v$ all non-zero should be considered. The solution of Eq. (8.10) yields a boson Green’s function which has a diffusive form at the critical point [293]
@@
This expression shows that the residual resistivity $\rho_{0}$ at $T=0$ is determined by the elastic scattering rate $1/\tau_{e}\sim{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}v^{2}}$, as in a disordered Fermi liquid. The inelastic processes lead to a frequency and temperature dependence which matches precisely with the observational form in Eq. (8.2). An important feature is that of the two processes contributing to the electron inelastic scattering rate $1/\tau_{\rm in}$ in Eq. (8.15), only one contributes to the inelastic transport rate $1/\tau_{\rm trans}$. The processes involving the spatially uniform interaction $g$ and the impurity potential $v$ cancel out in the computation of the conductivity from Fig. 38, and only those involving the spatially random interaction $g^{\prime}$ survive [122]. A consequence of this cancellation is that the constant $\widetilde{\alpha}$ in Eq. (8.5) approaches $\widetilde{\alpha}=\pi/2$ for the quasiparticle $m^{\ast}$ in the limit ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}g^{\prime}}\gg g$ [282], and decreases from this value as $g$ is increased [122].
To summarize, the conductivity of the theory $\mathcal{L}_{\Phi}$ yields the strange metal conductivity in Eq. (8.1), with $\rho_{0}\sim{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}v^{2}}$ and $A\sim{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}g^{\prime 2}}$. Note that the value of $g$ does not make a direct difference to the value of the linear-$T$ resistivity, although it does affect the marginal Fermi liquid behavior of the electron self energy, as noted in Eq. (8.15). It is also notable that the residual resistivity and linear-$T$ resistivity slope are determined by different sources of disorder: those in Eqs. (8.7) and (8.8) respectively. This distinction should be important in understanding trends in observations [294, 295]. In the quantum critical region, the approach outlined above [122], and the more complete quantum Monte Carlo studies [150] show that the slope $\widetilde{\alpha}$ of the transport scattering rate in Eq. (8.5) is largely independent of the strength of disorder.
@@
We note that a full numerical solution of Eq. (8.10) at $g=0$ has been presented by Li et al. [141], including results for the conductivity and the onset of superconductivity. Their results for the transport scattering time $\tau_{\rm trans}$ are shown in Fig. 39, along with the corresponding optical conductivity observations of Michon et al [270]. Both obey the Planckian scaling form in Eq. (8.2) reasonably well. The scaling is better in the experimental observations, and we suspect this is the due to the logarithmic violations in the theory, such as those in Eq. (8.16). It would be interesting to extend the study of optical conductivity to the quantum Monte Carlo approach [150] in future work.
@@ 9.5 Photoemission
This research was supported by NSF Grant DMR-2245246 and by the Simons Collaboration on Ultra-Quantum Matter which is a grant from the Simons Foundation (651440, S. S.). P.M.B. acknowledges support by the German National Academy of Sciences Leopoldina through Grant No. LPDS 2023-06 and the Gordon and Betty Moore Foundation’s EPiQS Initiative Grant GBMF8683. The Flatiron Institute is a division of the Simons Foundation.
Appendix A. Schrieffer-Wolff transformation from the ALM to the Hubbard model
Here we present an alternative derivation of the connection between the Hubbard model and the ALM, which is also illustrated in Fig. 7. We begin with a Hubbard model and augment it with ancillas which are decoupled from the Hubbard model. It is important that the new degrees of freedom have a trivial ground state with an energy gap, so that its extra excited states can be eliminated by a canonical transformation. For the ancillas, we choose a pair of qubits associated with every site of the square lattice, which we identify as $S=1/2$ spins ${\bm{S}}_{1{\bm{i}}}$ and ${\bm{S}}_{2{\bm{i}}}$. There is an antiferromagnetic exchange interaction $J_{\perp}$ between the two spins at each $\bm{i}$}, and so the ancillas form a bilayer square lattice antiferromagnet. We take $J_{\perp}$ to be large enough so that the ground state of the bilayer is smoothly connected to the ‘trivial’ rung singlet state. The decoupled ancilla Hamiltonian is therefore
$\mathcal{H}_{\rm ancilla-decoupled}=\mathcal{H}_{\rm Hubbard}+J_{\perp}\sum_{{\bm{i}}}{\bm{S}}_{1{\bm{i}}}\cdot{\bm{S}}_{2{\bm{i}}}+\ldots\,$, (.12)
where the ellipses represent exchange interactions within the ancilla layers. As indicated in Fig. 7, we proceed in a direction opposite to that of the paramagnon derivation. Starting from $\mathcal{H}_{\rm ALM}$, we perform a Schrieffer-Wolff transformation to map to $\mathcal{H}_{\rm ancilla-decoupled}$ [191]. This Schrieffer-Wolff transformation eliminates all the excited states of the ancilla spins in powers of $1/J_{\perp}$, and returns to the single band Hubbard model for the $c_{{\bm{i}}\alpha}$, with the value of $U$ shown below in Eq. (.16) and Fig. 7 [191]. At leading order in $1/J_{\perp}$, we can demonstrate this mapping by considering only a single value of $\bm{i}$}. Then at each $\bm{i}$}, the ancilla Hamiltonian is
$H_{1}=J_{\perp}{\bm{S}}_{1}\cdot{\bm{S}}_{2}-\mu c_{\alpha}^{\dagger}c_{\alpha}+J_{K}{\bm{S}}_{1}\cdot c_{\alpha}^{\dagger}\frac{{\bm{\sigma}}_{\alpha\beta}}{2}c_{\beta}$ (.13)
$E_{0}$ $=$ $-\frac{3J_{\perp}}{4},~~E_{2}=-\frac{3J_{\perp}}{4}-2\mu\,$,
$E_{1}$ $=$ $-\frac{3J_{\perp}}{4}-\mu-\frac{3J_{K}^{2}}{16J_{\perp}}+\mathcal{O}(1/J_{\perp}^{2})\,.$ (.14)
These states map onto those of a Hubbard model
$H_{2}=-\mu^{\prime}c_{\alpha}^{\dagger}c_{\alpha}+U(c_{\uparrow}^{\dagger}c_{\uparrow})(c_{\downarrow}^{\dagger}c_{\downarrow})$ (.15)
with parameters
$U=\frac{3J_{K}^{2}}{8J_{\perp}},\quad\mu^{\prime}=\mu+\frac{3J_{K}^{2}}{16J_{\perp}}\,.$ (.16)
Extending this mapping to all sites, we obtain the canonical (Schrieffer-Wolff) transformation connecting $\mathcal{H}_{\rm ALM}$ in Eq. (2.10) to $\mathcal{H}_{\rm ancilla-decoupled}$ in Eq. (.12).
Appendix B. Complete gauge theory of the ALM